If the length is 18 cm and the width is 18cm, what is the perimeter?
Found 2 solutions by robertb, mananth:
Answer by robertb(5830)

2*(18cm + 18cm) = 2*36 cm = 72 cm, the perimeter.

Answer by mananth(16949)

length = 18 cm
width = 18 cm
2*(L+W)=perimeter
2*(18+18)=72
Perimeter = 72 cm
...
m.ananth@hotmail.ca